OpenStudy (anonymous):

if the radius of the earth is roughly 3960 miles, how many times larger is the volume of the earth than the volume of a ping pong ball? A ping pong ball has a radius of 0.7441 inches

OpenStudy (anonymous):

\[Volume = \frac{ 4 }{ 3 } \times \pi \times r ^{3}\] Where r is the radius of the desired object. Calculate the volume of each sphere, then divide the results: earth/pingping = x times bigger